# ECC Analyzer: Extract Trading Signal from Earnings Conference Calls using Large Language Model for Stock Performance Prediction

### v1

@@ Abstract
In the realm of financial analytics, leveraging unstructured data, such as earnings conference calls (ECCs), to forecast stock performance is a critical challenge that has attracted both academics and investors. While previous studies have used deep learning-based models to obtain a general view of ECCs, they often fail to capture detailed, complex information. Our study introduces a novel framework: ECC Analyzer, combining Large Language Models (LLMs) and multi-modal techniques to extract richer, more predictive insights. The model begins by summarizing the transcript’s structure and analyzing the speakers’ mode and confidence level by detecting variations in tone and pitch for audio. This analysis helps investors form an overview perception of the ECCs. Moreover, this model uses the Retrieval-Augmented Generation (RAG) based methods to meticulously extract the focuses that have a significant impact on stock performance from an expert’s perspective, providing a more targeted analysis. The model goes a step further by enriching these extracted focuses with additional layers of analysis, such as sentiment and audio segment features. By integrating these insights, the ECC Analyzer performs multi-task predictions of stock performance, including volatility, value-at-risk (VaR), and return for different intervals. The results show that our model outperforms traditional analytic benchmarks, confirming the effectiveness of using advanced LLM techniques in financial analytics.
1 Introduction
The integration of structured data such as stock prices and financial ratios with unstructured data including financial filings and company news is increasingly essential in investment decision-making Fang and Zhang (2016); Roeder et al. (2022). This trend stems from the recognition that unstructured data can provide insights not fully captured by structured data alone. For instance, financial reports elucidate the broader implications of numerical data through discussions of managerial decisions and corporate strategies Wang and Hua (2014), while company news can shed light on public sentiment, emerging market trends, and market perceptions Kogan et al. (2009); Tetlock (2007). These nuanced insights are vital for a comprehensive analysis of complex market dynamics.
Advancements in Natural Language Processing (NLP) have significantly enhanced the ability to analyze unstructured data within the financial sector. Traditional NLP applications initially used bag-of-words models, which simplify text into isolated words, thus ignoring syntactical structure and sequence Zhang et al. (2010). Though effective in specific contexts like fraud detection Purda and Skillicorn (2015), these methods lack contextual understanding. More recent developments have introduced deep learning-based NLP algorithms, such as Embedding Language Models (ELMo) Mikolov et al. (2013) and Long Short-Term Memory Networks (LSTM) Hochreiter and Schmidhuber (1997), which provide a more effective capture of textual context. Nevertheless, these supervised learning-based methods are highly task-specific and have limited adaptability to generalization Singh et al. (2016).
The emergence of Large Language Models (LLMs) represents a paradigm shift in overcoming these limitations Brown et al. (2020); Touvron et al. (2023); Achiam et al. (2023). Equipped with expansive knowledge bases and sophisticated zero-shot learning capabilities, LLMs are capable of performing a diverse array of text-related tasks—ranging from summarization Zhang et al. (2024) and question-answering Wei et al. (2022) to sentiment analysis Zhang et al. (2023) and E-Commerce Jia et al. (2023) without the need for specialized training in specific tasks or domains Jia et al. (2022); Wang et al. (2023b); Yang et al. (2023); Gruver et al. (2024).
Drawing inspiration from recent advancements in LLMs, this work aims to explore the potential of LLMs to extract trading signals from earnings conference calls (ECCs) to improve predictions of stock performance. ECCs involve senior executives discussing quarterly results, providing a fertile ground for predicting stock movements through nuanced analysis of transcripts and audio. Despite the potential, existing research often overlooks finer details, lacks interpretability, or relies too heavily on sentiment analysis, leading to incomplete data interpretations.
Recent studies have increasingly leveraged domain-specific Large Language Models like FinBERT Liu et al. (2021), which is adapted from Google’s BERT algorithm for financial contexts. FinBERT evaluates the sentiment of ECCs by averaging the sentiment scores of each sentence, revealing a correlation between the sentiment and market reactions. Moreover, the integration of multimodal techniques has significantly improved the accuracy of financial risk predictions. For example, Qin and Yang (2019) and Yang et al. (2020) utilize both textual and auditory data to generate embeddings that encapsulate semantic and auditory features. These models have proven effective in predicting market volatility at various intervals, demonstrating the potential of combining textual analysis with auditory data processing. While these studies may get the descent performance of prediction, another concern with using deep learning-based or LLM techniques is the challenge of explaining how the model arrives at its results. Several interpretability models have been proposed to explain the model’s decision reasons.  Wang et al. (2023a); Tenney et al. (2020).
Existing research, however, reveals significant gaps that warrant further exploration: 1) the previous studies directly input entire texts or audio files into models, potentially missing important details and lacking interpretability, especially regarding which earnings call topics influence predictions; 2) While some studies focus on sentiment extraction from ECCs, they capture only a fraction of the available information, leading to potentially incomplete interpretations. The findings suggest that sentiment analysis alone offers limited explanatory power for predicting stock movements, highlighting the need for broader data utilization; 3) Additionally, integrating large language models (LLMs) into financial analysis while ensuring that investors understand the reasoning behind model outputs poses a distinct challenge, prompting ongoing research.
Given these gaps, in designing a framework to extract trading signals from ECCs to predict financial performance, this paper is interested in exploring the following research questions (RQs):
  • RQ1: How can large language models be used to provide investors with a more comprehensive understanding of a company’s financial health and strategic direction?
  • RQ2: Can a more comprehensive analysis provide additional predictive capability for stock performance?
  • RQ3: Can LLMs be employed to generate interpretable content that aids investors in understanding the decision-making process?
To address the above RQs, this paper introduces the ECC Analyzer, a novel framework utilizing LLMs for in-depth analysis of ECC data. The framework initially provides an understanding of ECCs by segmenting transcripts into themes like financial performance and corporation project discussions. It summarizes these segments, distilling the essence of each thematic chunk, and then combines these summaries into a comprehensive overview. This hierarchical summarization enables stakeholders to grasp complex documents’ main themes and insights. Regarding audio, the model analyzes speech features such as tone, pitch, and intensity to gauge the speaker’s confidence level.
Furthermore, the ECC Analyzer simulates how investors examine key indicators and infer future market behavior. We begin by creating a database (focus) of key indicators with finance experts, such as financial metrics, management changes, operational costs, and strategic plans. Using the constructed database, the ECC Analyzer employs retrieval-augmented generation (RAG) to systematically examine ECCs and pinpoint factors critical to investment decisions. After thoroughly extracting and analyzing the ECC, the results are integrated with raw ECC data to conduct multi-task learning: predicting volatility as well as Value at Risk (VaR) and return for different time intervals (3, 7, 15, and 30 days).
By utilizing RAG, our method improves model interpretability by linking specific earnings call topics directly to stock performance, enhancing both the investor’s understanding and the explainability of the analysis. To validate our approach, we benchmarked our model against traditional methods, showing significant improvements in accuracy and predictive power. These results highlight our model’s utility for financial analysts and investors in making informed decisions based on detailed data analysis.
2 Related Work
This paper integrates concepts from stock performance prediction and advancements in large language models (LLMs). In this section, we offer a concise review of key research in both areas relevant to our study.
2.1 Stock Performance Prediction
@@
The use of NLP techniques to analyze unstructured data for predicting stock performance has attracted significant academic attention. A foundational study by Kogan et al. (2009) shows that simple bag-of-words features from annual reports when combined with historical volatility, can outperform models based solely on historical data. Subsequent research, such as that by Wang and Hua (2014); Rekabsaz et al. (2017); Theil et al. (2018), proposed various document representation methods to predict stock price volatility. Drawing on multimodal technologies,  Qin and Yang (2019) explored how audio features—such as tone, emotion, and speech rate—enhance stock movement predictions when combined with text analysis. Following by this, Yang et al. (2020) further extends the idea of using multimodal data to improve risk prediction performance in multi-task learning, and the authors’ experiments show that predicting multiple tasks at the same time can help the model further improve prediction performance. However, the aforementioned studies primarily input ECC data directly into models for prediction without conducting a thorough analysis of the ECC content.
2.2 Large Language Models in Finance
@@
Numerous studies have explored the applications of LLMs in the financial sector.  Li et al. (2023) explore how LLMs have been adeptly applied to summarize and abstract complex financial documents such as 10-K, and 10-Q filings. Yang et al. (2023); Yu et al. (2023) explores the usage of LLMs in mining media news for trading recommendations, showcasing the models’ ability to discern subtle market indicators and sentiments. In the domain of customer service, the implementation of LLM-powered chatbots is spotlighted for offering context-aware interactions, serving as both assistants and consultants Lakhani (2023); Subagja et al. (2023); Soni (2023)Abdaljalil and Bouamor (2021); Zmandar et al. (2021) explore the nuanced task of extracting financial and legal items from lengthy text documents, such as financial regulations and comprehensive policy manuals. However, these existing studies predominantly focus on tasks like financial text summarization, question-answering (Q&A), and stock movement prediction (binary classification), with a notable gap in the application of LLMs for comprehensive stock performance prediction.
Figure 1: This figure illustrates the ECCs Analyzer’s Framework. The model accepts multimodal inputs: ECC Audio and Transcript. The second area visualizes the model’s pipeline to encode text and audio sources and illustrates how LLMs are applied for information analysis. This model analyzes text-audio pairs, summarizes ECCs, and extracts focal points of interest from a financial expert’s perspective. The third area describes how the model consolidates outputs from both embeddings and LLM analysis for use in subsequent stages. The model will perform multi-task learning: our ECCs Analyzer will predict the Volatility of different terms and VaR in the meantime.
3 Methodology
ECC Analyzer, illustrated in Figure (1) aims to comprehensively understand the multi-data types present in earnings conference calls, including both text and audio components. (3.1) audio encoding (3.2) transcript encoding. Furthermore, it focuses on extracting trading signals from the analyzed data to predict stock performance (3.3) earnings conference call focuses extraction. The model includes a component to optimally integrate different data sources (3.4) additive multi-model fusion. The model performs multi-task prediction: to predict the the volatility across different terms and the value at risk (VaR) at the same time (3.5). It also demonstrates strong predictive ability in forecasting stock returns.
Figure 2: This figure visualizes the mechanism of the earnings conference call analyzer. This framework takes both earnings conference calls’ audio and transcript as input. It starts by hierarchically summarizing the content of ECCs. Further, this model simulates the process through which an investor analyzes an ECC. It examines and extracts focuses on interest, such as financial indicators and business events, from a database designed by financial experts. Subsequently, this model calculates the sentiment for these focal points and extracts audio features, such as changes in tone and pitch, from the corresponding audio segments. All these analyses are combined to form a comprehensive encoder for further use.
3.1 Audio Encoding
@@
Audio pre-trained models have achieved performing results in various downstream tasks Pons and Serra (2019); Cramer et al. (2019); Koh and Dubnov (2021); Wang et al. (2021). We aim to leverage advanced audio pre-trained models like Wav2vec2, a transformer-based Large Language Model recognized for its effectiveness in processing raw audio Baevski et al. (2020), to extract audio embeddings. After that, we employ a Multi-Head Self-Attention (MHSA) mechanism to distill specific audio features. This method is vital for integrating these features with other data modalities, facilitating a more detailed and comprehensive analysis.
@@
To describe the Audio Encoding in more detail, we let the raw audio input data be represented by $A_{c}=\{a_{c}^{1},a_{c}^{2},\ldots,a_{c}^{n}\}$ where $a_{c}^{i}$ represents the $i^{th}$ audio frame in one data sample. Each audio frame will be converted into a vector representation: $e_{ac}^{i}=\text{Wav2Vec2}(a_{c}^{i})$. Therefore, we obtain the audio embeddings $E_{ac}=\{e_{ac}^{1},e_{ac}^{2},\ldots,e_{ac}^{n}\}$ which have dimensions of 520 $\times$ 512, representing the maximum number of audio files across companies and the transform dimensions for a single audio frame, respectively. Audio files with fewer than 520 frames ($n<520$) are zero-padded for consistent matrix size.
@@
$\text{Multihead}=\text{Concat}\text{(head}_{1},....,\text{head}_{h})W^{o}$ (1)
$\text{head}_{i}=\text{Attention}(QW_{i}^{Q},KW_{i}^{K},VW_{i}^{V})$ (2)
@@
$\text{Attention}(Q,K,V)=softmax(\frac{KQ^{T}}{\sqrt{d_{k}}})V$ (3)
@@ 3.2 Transcript Encoding
The transcription process mirrors Audio Encoding, using SimCSE Gao et al. (2021)to extract sentence-level vector representations from earnings conference transcripts. SimCSE is a Siamese neural network architecture that learns to embed pairs of sentences into a shared space where similar sentences are mapped close together and dissimilar sentences are mapped far apart. The raw transcripts are represented as $T_{c}\ =\{t_{c}^{1},a_{c}^{2},\ldots,t_{c}^{n}\}$, with each sentence $t_{c}^{i}$ represents the $i^{th}$ transformed into a vector representation: $e_{tc}^{i}=\text{SimCSE}(t_{c}^{i}).$
@@
We obtain the corresponding text embeddings given by $E_{tc}=\{e_{tc}^{1},e_{tc}^{2},\ldots,e_{tc}^{n}\}$ with size 520 $\times$ 768, where 520 is the maximum number of sentences amongst all data samples and 768 is the dimension of the output of SimCSE. Earnings conference calls with less than 520 sentences ($n<520$) have been zero-padded for uniformity in input matrix size. Same with (1)-(4), the MHSA is applied to $E_{tc}$ to get $T_{tc}=\{t_{tc}^{1},t_{tc}^{2},\ldots,t_{tc}^{n}\}$ with dimension 520 $\times$ 768. Then,$T_{tc}$ is subjected to the average pooling layer to produce $T_{t}$, where $T_{t}$ denotes the resultant extracted textual feature of size 768. This streamlined process effectively captures the textual nuances required for in-depth analysis.
3.3 Earnings Conference Call Focuses Extraction And Analysis
To obtain deep insights from an Earnings Conference Call on how it might influence future market performance, our approach encompasses several steps: (1) summarize ECCs, (2) extract investors’ focus information (3) calculate focus sentiment and (4) extract audio’s features of corresponding focus. Refer to the Figure 2 for additional details.
(1) Summarize ECCs. To accommodate the token limitations of LLMs, we start by dividing the entire document into several chunks and summarizing each one. We then summary these summaries to create a comprehensive overview of the entire document. This two-tiered approach ensures that the summary capture both the detailed and overall information. In further, we will use LLM to get the embedding with a size 1024: $T_{s}=\text{LLM}(summary+chunk\ summaries)$
(2) Extract the investor’s focus using RAG. In this step, we attempt to use LLM to examine and extract focuses of interest to investors. Therefore, we start by identifying the core topics that analysts typically focus on. We consult with four finance experts, each possessing extensive experience in financial analysis. These experts are tasked with meticulously reviewing earnings conference calls across various sectors to pinpoint and summarize the key topics that are frequently discussed and also hold the greatest interest for investors. Their analysis aims to identify overlaps where common discussion topics align closely with investor concerns, highlighting areas of particular significance and interest in the investment community. Once we have established this focused database, our goal is to locate where these topics are mentioned during the earnings calls and systematically extract these segments. However, accurately extracting this information from the dense narrative of earnings conference calls can pose a challenge. Because LLM searches are based on similarity, comparing a single-word topic to an entire document presents a scale issue, and a single word may lack the necessary context, thus diminishing the extraction capability. To enhance the precision of information extraction, we formulate multiple questions related to each topic but phrased differently, ensuring a broad coverage and increasing the likelihood of accurately pinpointing relevant sections. For instance, regarding the dividends topic, we pose questions such as: (1) "Have there been any changes in the stock dividends, and if so, at what rate have the dividends increased or decreased?" (2) Does the company Board expect to increase the stock dividends in the next future? (3) How does the dividend yield compare to the peers?
After locating the segments where these topics are discussed, we extract the corresponding paragraphs. These paragraphs, however, may still contain irrelevant text such as preceding discussions. To address this, we use a contextual compression method from LLM, adept at distilling the paragraph down to its most relevant sentences, effectively obtaining the crucial data from the irrelevant text.
(3) Calculate focus sentiment. Sentiment is a strong indicator for reflecting market perception. We determined to calculate the sentiment score based on the focus. Because this offers a more targeted and insightful view, which may be overlooked in a broader context. It provides a clearer picture of how specific developments or concerns are influencing investor behavior and market movements.
(4) Extract audio’s features using RAG. Once we identify the focal points of interest, we aim to enrich our analysis by locating and examining the audio segments that correspond to these specific topics. By analyzing the audio features related to a particular focus, such as the company’s current projects, we can delve deeper into the emotional nuances and confidence levels exhibited by the speakers. This approach allows us to capture subtle cues in tone, pace, and emphasis that might indicate underlying sentiments or confidence about the discussed topics. The goal is to provide a more layered understanding of how speakers convey their messages and the potential impacts these emotional expressions have on the listeners’ perceptions and reactions. We utilize Praat Boersma and Van Heuven (2001) to extract vocal features, such as pitch, intensity, jitter, HNR(Harmonic to Noise Ratio) and etc, from ECC audio files.
After obtaining the four key components, we then explore effective methods to merge these elements into a cohesive analysis. In further, we will use LLM to get the embedding with size 1024: $T_{f}=\text{LLM}(focus\ analysis+focus\ sentiment+focus\ audio\ features)$
3.4 Additive Multi-modal Fusion

### v2

# ECC Analyzer: Extract Trading Signal from Earnings Conference Calls using Large Language Model for Stock Volatility Prediction

@@ Abstract.
In the realm of financial analytics, leveraging unstructured data, such as earnings conference calls (ECCs), to forecast stock volatility is a critical challenge that has attracted both academics and investors. While previous studies have used multimodal deep learning-based models to obtain a general view of ECCs for volatility predicting, they often fail to capture detailed, complex information. Our research introduces a novel framework: ECC Analyzer, which utilizes large language models (LLMs) to extract richer, more predictive content from ECCs to aid the model’s prediction performance. We use the pre-trained large models to extract textual and audio features from ECCs and implement a hierarchical information extraction strategy to extract more fine-grained information. This strategy first extracts paragraph-level general information by summarizing the text and then extracts fine-grained focus sentences using Retrieval-Augmented Generation (RAG). These features are then fused through multimodal feature fusion to perform volatility prediction. Experimental results demonstrate that our model outperforms traditional analytical benchmarks, confirming the effectiveness of advanced LLM techniques in financial analysis.
Large Language Model, Earnings Conference Call Analysis, Volatility forecasting, Retrieval-Augmented Generation
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2024; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Natural language processingccs: Information systems Multimedia information systems
1. Introduction
Predicting the stock volatility over a certain period is a crucial task in financial analysis, aiding capital market participants in making better investment decisions. As such, developing effective techniques for predicting stock volatility has become increasingly important among academics and the financial industry. Previous studies in economics have shown that stock volatility can be predicted using publicly available information (Dumas et al., 2009). Substantial research in finance and computational linguistics has made significant progress in predicting stock volatility from various textual sources, such as company disclosed reports (Kogan et al., 2009; Hoberg and Phillips, 2016; Rekabsaz et al., 2017), the transcripts of earnings conference calls (Kimbrough, 2005; Wang and Hua, 2014) and social media (Bollen et al., 2011; Ding et al., 2015; Oliveira et al., 2017).
Recently, advancements in multimodal learning have enabled the use of more unstructured multimedia data, such as audio recordings of earning conference calls (Qin and Yang, 2019; Li et al., 2020), Merge & Acquisition calls (Sawhney et al., 2021) and the video of CEO’s speech (Rajandran, 2021), for stock volatility prediction. Multimodal learning is particularly valuable in this context as it allows for the integration of diverse data sources, providing richer, more nuanced insights into market dynamics and sentiment. This comprehensive analysis is essential for understanding complex market behaviors (Qin and Yang, 2019). Our study focuses on multimodal earning conference calls (ECCs) data for two primary reasons: 1) transcripts and audio recordings are publicly available, and 2) ECCs are often associated with high volatility primarily due to the market reaction to the earnings announcement (Foster et al., 1984).
Existing multimodal approaches to volatility prediction using ECCs typically extract features from speech and text separately. Subsequently, commonly used Natural Language Processing (NLP) models, such as LSTM or Transformer-based models, are employed to jointly model the extracted speech and text features, ultimately performing volatility prediction (Qin and Yang, 2019; Yang et al., 2020; Wang et al., 2024). While these methods have shown that the multimodal approach can extract complementary information from multiple modalities, enhancing financial modeling performance and model robustness, several challenges remain unaddressed: firstly, previous work directly feeds features extracted from text and audio into the model, potentially missing important contextual details; secondly, these approaches assign equal importance to all sentences and audio segments which hard to reflect the impact of important information in the ECCs on the prediction.
Figure 1. illustrates the ECC Analyzer Framework. The proposed method accepts multimodal inputs: audio record and transcript. The upper part of the box illustrates the feature extraction process for both the audio and text of the data. We use the pre-trained large models to generate embeddings from the audio and text, followed by a transformer encoder to extract the corresponding features. The lower part of the box represents a deeper analysis of the ECC. Here, the text is divided into chunks, which are then summarized into paragraphs using an LLM. Key sentences are extracted via RAG and converted into text features through text embedding. These text features are then fused with the features extracted from the upper part of the box to make the final volatility prediction.
In response to the aforementioned limitations and inspired by the significant performance improvements offered by large language models (LLMs) in various downstream NLP tasks, this paper introduces the ECC Analyzer — a novel framework that leverages LLMs for in-depth analysis of ECC data to enhance volatility prediction performance. The proposed ECC Analyzer employs a hierarchical information extraction strategy: 1) advanced pre-trained large models are used to extract embeddings from both audio and text, capturing the overall information content of ECCs.; 2) LLMs summarize ECC transcripts at the paragraph level, distilling important information from the transcript; 3) In collaboration with financial experts, we designed the “Question Bank” which contains a series of questions about ECC content that are of interest to investors. These questions are used as queries to extract fine-grained sentences containing key information from ECCs using Retrieval-Augmented Generation (RAG). We use these questions as queries and extract fine-grained sentences containing key information from ECC by RAG to improve the prediction performance of the model. Then the summarized text with extracted key focus sentences is transformed into text embedding and combined with the extracted audio and text features for volatility prediction.
Our extensive experiments on the real-world S&P 500 ECCs dataset demonstrate clear and significant improvements in prediction accuracy attributable to our proposed approach. We claim that the contribution is two-fold: 1) Our framework innovatively integrates multi-modal information with fine-grained details extracted by Large Language Models (LLMs) to effectively distill and utilize rich information from financial documents, such as ECCs, for prediction tasks. 2) Our pipeline demonstrates a $27.7\%$ reduction in average Mean Squared Error (MSE) compared to the current state-of-the-art (SOTA) model. In detail, The results show that our method achieves substantial performance improvements in forecasting short-term volatility (3-day, 7-day) compared to current SOTA methods. The results for medium-term volatility (15 days) and long-term volatility (30 days) are comparable to the best existing methods.
2. Related Work
2.1. Stock Volatility Prediction
@@
Volatility forecasting has long been of interest to researchers due to its practical applications. Conventional forecasting approaches rely on historical stock prices and use continuous and discrete time series models (Kristjanpoller et al., 2014; Cox and Ross, 1976; Heston, 1993; Hsu et al., 2008; Franses and Van Dijk, 1996; Kim and Won, 2018). Recently, the use of NLP techniques to analyze unstructured data for predicting stock performance has attracted significant academic attention. A foundational study by (Kogan et al., 2009) shows that simple bag-of-words features from annual reports when combined with historical volatility, can outperform models based solely on historical data. Subsequent research, such as that by (Wang and Hua, 2014; Rekabsaz et al., 2017; Theil et al., 2018), proposed various document representation methods to predict stock price volatility. Drawing on multimodal technologies,  (Qin and Yang, 2019) explored how audio features—such as tone, emotion, and speech rate—enhance stock movement predictions when combined with text analysis. Following by this, (Yang et al., 2020) further extends the idea of using multimodal data to improve risk prediction performance in multi-task learning, and the authors’ experiments show that predicting multiple tasks at the same time can help the model further improve prediction performance. (Wang et al., 2024) addressed the reduction of gender bias in ECC predictions through adversarial training. However, the aforementioned studies primarily input ECC data directly into models for prediction without conducting a thorough analysis of the ECC content.
2.2. Large Language Models in Financial Application
@@
Numerous studies have explored the applications of LLMs in the financial sector.  (Li et al., 2023) explore how LLMs have been adeptly applied to summarize and abstract complex financial documents such as 10-K, and 10-Q filings. The FinBen (Xie et al., 2024) provides the benchmark performance of LLM for each task in the financial domain. FinGPT (Yang et al., 2023) and FinMem (Yu et al., 2023) explore the usage of LLMs in mining media news for trading recommendations, showcasing the models’ ability to discern subtle market indicators and sentiments. In the domain of customer service, the implementation of LLM-powered chatbots is spotlighted for offering context-aware interactions, serving as both assistants and consultants (Lakhani, 2023; Subagja et al., 2023; Soni, 2023)(Abdaljalil and Bouamor, 2021; Zmandar et al., 2021) explore the nuanced task of extracting financial and legal items from lengthy text documents, such as financial regulations and comprehensive policy manuals. However, these existing studies predominantly focus on tasks like financial text summarization, question-answering (Q&A), and stock movement prediction (binary classification), with a notable gap in the application of LLMs for comprehensive stock volatility prediction.
3. Problem Formulation
Volatility can be expressed as the natural log of the standard deviation of return prices $r$ in a window of ${\tau}$ days. We calculate the 3, 7, 15, and 30 trading days volatility using the equation:
(1) $v_{[d-\tau,d]}=\ln\left(\left(\frac{\sum_{i=0}^{\tau}(r_{d-i}-\bar{r})^{2}}{\tau}\right)^{\frac{1}{2}}\right)$
The notations used across the paper are discussed herein. Let $s\in S$ denote a stock, $c\in C$ be an earnings call for stock s. For each stock, there exist multiple earnings calls c that are held periodically. Each call c can be segmented into a set of $a_{c}^{i}\in A_{c}$ audio clips, and corresponding $t_{c}^{i}\in T_{c}$ text sentences for $i\in[1,N]$, where is the $N$ is the maximum number of audio clips in a call. For each stock, there exists a daily return denoted by $r_{i}=\ \frac{p_{i}-p_{i-1}}{p_{i-1}}$, where $p_{i}$ is the adjusted close price at the end of the day and r̄ is the average return over a period of ${\tau}$ days. We aims to develop a predictive regression function $f(c)\rightarrow v_{[d-\tau,d]}$
For the evaluation, following (Qin and Yang, 2019), we assess the accuracy of volatility predictions by comparing the predicted values $y_{i}$ with the labeled volatility values $\hat{y_{i}}$. We use Mean Squared Error (MSE) as the evaluation metric:
(2) $MSE=\frac{\sum_{i}(y_{i}-\hat{y_{i}})^{2}}{n}$
This metric quantifies the average squared difference between the predicted and actual volatility values, providing a measure of prediction accuracy. Additionally, for constructing volatility labels, we collected stock price information via the Yahoo Finance API [^1] https://finance.yahoo.com/ and used equation 1 to calculate the labels for days 3, 7, 15, and 30.
4. Our proposed framework
ECC Analyzer (in Figure 1) aims to comprehensively understand the multi-data types present in earnings conference calls, including both text and audio components. In this section, we explain our methodology in 4 parts: (4.1) audio encoding, (4.2) transcript encoding,(4.3) fine-grained information extraction from the transcript by using LLM, and (4.4) multi-model fusion and model training.
4.1. Audio Encoding
@@
Audio pre-trained models have achieved performing results in various downstream tasks (Pons and Serra, 2019; Cramer et al., 2019; Koh and Dubnov, 2021; Wang et al., 2021). We aim to leverage advanced audio pre-trained models like Wav2vec2, a transformer-based Large Language Model recognized for its effectiveness in processing raw audio (Baevski et al., 2020), to extract audio embeddings. After that, we employ a Multi-Head Self-Attention (MHSA) mechanism to distill specific audio features. This method is vital for integrating these features with other data modalities, facilitating a more detailed and comprehensive analysis.
Figure 2. visualizes the process of fine-grained information extraction from ECC transcript.
@@
The raw audio input data be represented by $A_{c}=\{a_{c}^{1},a_{c}^{2},\ldots,a_{c}^{n}\}$ where $a_{c}^{i}$ represents the $i^{th}$ audio frame in one data sample. Each audio frame will be converted into a vector representation: $e_{ac}^{i}=\text{Wav2Vec2}(a_{c}^{i})$. Therefore, we obtain the audio embeddings $E_{ac}=\{e_{ac}^{1},e_{ac}^{2},\ldots,e_{ac}^{n}\}$ which have dimensions of 520 $\times$ 512, representing the maximum number of audio files across companies and the transform dimensions for a single audio frame, respectively. Audio files with fewer than 520 frames ($n<520$) are zero-padded for consistent matrix size.
@@
(3) $\text{Multihead}=\text{Concat}\text{(head}_{1},....,\text{head}_{h})W^{o}$
(4) $\text{head}_{i}=\text{Attention}(QW_{i}^{Q},KW_{i}^{K},VW_{i}^{V})$
@@
(5) $\text{Attention}(Q,K,V)=softmax(\frac{KQ^{T}}{\sqrt{d_{k}}})V$
@@ 4.2. Transcript Encoding
The transcript encoding process mirrors Audio Encoding, using SimCSE (Gao et al., 2021) to extract sentence-level vector representations from earnings conference transcripts. SimCSE is a Siamese neural network architecture that learns to embed pairs of sentences into a shared space where similar sentences are mapped close together and dissimilar sentences are mapped far apart. The raw transcripts are represented as $T_{c}\ =\{t_{c}^{1},a_{c}^{2},\ldots,t_{c}^{n}\}$, with each sentence $t_{c}^{i}$ represents the $i^{th}$ transformed into a vector representation: $e_{tc}^{i}=\text{SimCSE}(t_{c}^{i}).$
@@
We obtain the corresponding text embeddings given by $E_{tc}=\{e_{tc}^{1},e_{tc}^{2},\ldots,e_{tc}^{n}\}$ with size 520 $\times$ 768, where 520 is the maximum number of sentences amongst all data samples and 768 is the dimension of the output of SimCSE. Earnings conference calls with less than 520 sentences ($n<520$) have been zero-padded for uniformity in input matrix size. Same with (1)-(4), the MHSA is applied to $E_{tc}$ to get $T_{tc}=\{t_{tc}^{1},t_{tc}^{2},\ldots,t_{tc}^{n}\}$ with dimension 520 $\times$ 768. Then,$T_{tc}$ is subjected to the average pooling layer to produce $T_{t}$, where $T_{t}$ denotes the resultant extracted textual feature of size 768.
4.3. Fine-grained Information Extraction from the Transcript by Using LLM
To obtain deep insights from an Earnings Conference Call on how it might predict future market performance, our approach encompasses two parts: (1) summarize the transcript, and (2) important sentences extracted by using RAG. We show this process in Figure 2. We list all Prompts in Appendix A.1.
4.3.1. Summarize the transcript
In order to effectively summarize key information in lengthy ECC records, we employ a hierarchical summarization strategy. First, the entire document is divided into chunks, and then we use LLM to summarize each chunk individually. These individual summaries are then further summarized through LLM, resulting in a comprehensive summarization paragraph of the entire document. This two-layer approach ensures that both detailed and aggregate information is captured. In further, we use the OpenAI ‘text-embedding-3-small’ text embedding model to generate embeddings $T_{s}$ with size 1024 for summarized paragraph:
(7) $T_{s}=\text{Embedding}(summary+chunk\ summaries)$
4.3.2. Important sentences extracted by using RAG
In this step, we aim to extract the most critical sentences containing relevant information from the entire transcript. To achieve this, we first convert the processed chunks into vector representations using the embedding model. Then, with input from financial experts, we design a list of questions about the ECC, which are stored in a ’Question Bank’ as a query list (see Appendix  A.2). For each question in the ’Question Bank’, we retrieve the relevant chunks from the vectors individually. Next, we use the context compressor, in combination with the LLM chat, to quiz each extracted chunk and query on its relevance. This process filters out only the sentences useful for answering the questions. Finally, we synthesize the query-selected sentences into a coherent prompt, to which the large language model will generate responses as the final answers to the queries. These generated answers are considered as the extracted important sentences. We then concatenate all the generated sentences and feed them into a text embedding model to generate the text embedding $T_{f}$ with size 1024:
(8) $T_{f}=\text{Embedding}(\text{Concatenated}[selected\ sentences;])$
4.4. Multimodal Fusion and Prediction